a polygon has n sides.two of its angles are right-angles.each of the remaining angles is equal to equal to 150 degrees.calculate n.
Answer by J2R2R(94)

For a polygon of n sides, the sum of all angles is 180n – 360 degrees

We are given that two sides are 90 degrees and the remainder are 150 degrees

Therefore 180n – 360 = 150(n – 2) + 90 x 2 = 150n – 300 + 180

180n = 150n + 360 + 180 – 300

30n = 240

n = 8

Therefore it is an octagon with six angles 150 degrees and two right angles.